### 17 Dec 2023 v1

@@ 2.0.1 Fine-Tuned Models
Named entity recognition (NER) and RE from publications in ML and related fields have been tackled by SciERC [19] and subsequently SciREX [13]. The scope considered are methods, tasks, data sets, and evaluation metrics. Proposed methods for the task utilize BiLSTMs, BERT and SciBERT [5]. With both approaches, there is a partial overlap in entity types to our task, as we also extract methods and data sets. The key difference arises though the nature of parameters and values we relate them to, and the challenges in extracting those as briefly laid out in the introduction.
@@ 4.2 LLM
We find (1) to lead to problems with errors propagation along steps, and with (2) and (4) we frequently see alterations in the reproduced text. Accordingly, we use prompt type (3) for our approach—specifying a data serialization format in the prompt. While existing work uses the JSON format for this [11], we use YAML, as it is less prone to “delimiter collision” problems due to its minimal requirements for structural characters.111111See https://yaml.org/spec/1.2.2/. In doing so, we expect to avoid problems with LLM output not being parsable. Our overall LLM approach looks as follows.
@@
Few-shot: Our few shot approach makes the following adjustments to the method described above. Prompts additionally include a component [examples], which are valid input output pairs sampled by their similarity to the input text. Specifically, for an input text from a document X, we sample the 5 most similar paragraphs from all ground truth documents excluding X. As these examples can be confused with the input text, we place it after the examples, resulting in the structure [task][format][examples][input text][completion prefix].
LLMs reaching a sufficient context size for a few-shot approach to our task are a recent development. We can therefore additionally make use of other recently added capabilities. Specifically, we make use of generation constrains via a gBNF grammar121212See https://github.com/ggerganov/llama.cpp/pull/1773. to enforce LLM output according to our data scheme, allowing us to mitigate parsing errors.
@@ 5.1 Fine-Tuned Models
We use PL-Marker, the currently best performing model on SciERC, as our baseline. Models are trained and evaluated using 5-fold cross validation (3 folds training, 1 dev, 1 test). We train the ER component of PL-Marker as done in [33], using scibert-scivocab-uncased as the encoder, Adam as the optimizer, a learning rate of 2e-5, and for 50 epochs. The PL-Marker RE component is trained using bert-base-uncased, Adam, a learning rate of 2e-5, and for 10 epochs. Our own RE component also uses bert-base-uncased, Adam as the optimizer, and is trained with a learning rate of 1e-3 for 90 epochs. Models are trained and evaluated on a single GeForce RTX 3090.
@@ 5.1.2 Analysis
Token level ER performance across entity classes (none, artifact, parameter, value, context) is at 98.5%, 77.8%, 47.9%, 84.4%, 0% . That is, the model does predict contexts and struggles with parameters, but artifacts and values are predicted reliably. For our RE model, we observe that value-parameter relations are more reliably predicted than parameter-artifact relations.
@@ 5.2 LLMs
The models are run as follows. GPT-3.5 is accessed through its official API. All open models are run on a high performance compute cluster. Vicuna and WizardLM are run on nodes with NVIDIA Tesla V100. GALACTICA, Falcon, and Vicuna are run on nodes with NVIDIA A100.
@@ 5.2.2 Analysis
In Figure 6 we show an analysis of the steps leading up to model prediction. Focussing first on the zero-shot models (upper five) we observe the following across the four plots from left to right. (a) For three of five models, prompting for YAML leads to fewer parsing errors. (b) Unwanted text around the extracted data is generated more/less by two models each. (c) Hallucinated entities and (d) out of scope entities appear overall slightly more often for in YAML compared to JSON. For our few shot approach (bottom model), we see that the use of a grammar (a, b) prevents all output format issues. Furthermore (c) hallucinated entities are reduced. (d) Out of scope entities can not be evaluated, because our in-context examples lead to frequent omission of type information in the output.
@@ 6 Discussion
Our overall results, with a top performance of 79% for ER and 39% for RE, show that HyperPIE can be accomplished to a degree that yields sound results, but challenges still remain. Our novel data set enables further development of approaches from hereon. Our IE results on large-scale unannotated data give an indication of possible downstream analyses and applications. Here we see large potential for reproducibility research, faceted search, and recommendation.
Our LLM evaluation shows that for IE tasks dealing with complex information, the choice of text based data serialization format can have a considerable impact on performance, even when using grammar based generation constrains. Additionally, we can see that in-context learning enabled by larger context sizes, as well as grammars are an effective method to improve IE performance.
Limitations: Our work considers HyperPIE from text. This is sensible for a focussed approach, but downstream applications could furthermore benefit from composite pipelines also targeting extraction from tables, source code, etc. Our work does not test transferability of methods to domains outside of ML related fields. It would require domain expertise to find useful definitions for hyperparameters in each respective domain. Lastly, our data and experiments unfortunately are limited to English text only and do not cover other languages.

### 10 Jan 2024 v2

@@ 2.0.1 Fine-Tuned Models
Named entity recognition (NER) and RE from publications in ML and related fields have been tackled by SciERC [19] and subsequently SciREX [13]. The entity types considered are methods, tasks, data sets, and evaluation metrics. Proposed methods for the task utilize BiLSTMs, BERT and SciBERT [5]. With both approaches, there is a partial overlap in entity types to our task, as we also extract methods and data sets. The key difference arises though the parameter and value entities we cover, which are a challenge in part due to their varied forms of notation (e.g. / alpha, or / / 1e-3).
@@ 4.2 LLM
We find (1) to lead to problems with errors propagation along steps. With (2) and (4) we frequently see alterations in the reproduced text. Accordingly, we use prompt type (3) for our approach—specifying a data serialization format in the prompt. While existing work uses the JSON format for this [11], we use YAML, as it is less prone to “delimiter collision” problems due to its minimal requirements for structural characters.111111See https://yaml.org/spec/1.2.2/. In doing so, we expect to avoid problems with LLM output not being parsable. Our overall LLM approach looks as follows.
4.2.1 Zero-shot
@@
4.2.2 Few-shot
Our few-shot approach makes the following adjustments to the method described above. Prompts additionally include a component [examples], which are valid input output pairs sampled by their cosine similarity to the input text. Specifically, for an input text from a document X, we sample the five most similar paragraphs from all ground truth documents excluding X. As these examples can be confused with the input text, we re-position the input text to appear after the examples. The resulting prompt structure we use for our few-shot approach is as follows: [task][format][examples][input text][completion prefix].
LLMs reaching a sufficient context size for a few-shot approach to our task are a recent development. We can therefore additionally make use of other recently added capabilities. Specifically, we make use of generation constrains via a gBNF grammar131313See https://github.com/ggerganov/llama.cpp/pull/1773. to enforce LLM output according to our data scheme, allowing us to mitigate parsing errors.
@@ 5.1 Fine-Tuned Models
We use PL-Marker, the currently best performing model on SciERC, as our baseline. Models are trained and evaluated using 5-fold cross validation (3 folds training, 1 dev, 1 test). We train the ER component of PL-Marker as done in [33], using scibert-scivocab-uncased as the encoder, Adam as the optimizer, a learning rate of 2e-5, and 50 training epochs. Regarding the two RE components we compare, the PL-Marker RE component is trained using bert-base-uncased, Adam, a learning rate of 2e-5, and 10 training epochs. Our own RE component also uses bert-base-uncased, Adam as the optimizer, and is trained with a learning rate of 1e-3 for 90 epochs.141414The two RE models we compare require different learning rates and number of training epochs, because their architecture varies significantly. The models are trained and evaluated on a server with a single GeForce RTX 3090 (24 GB).
@@ 5.1.2 Analysis
Token level ER performance across entity classes (none, artifact, parameter, value, context) is at 98.5%, 77.8%, 47.9%, 84.4%, 0% . That is, the model does not predict contexts and struggles with parameters, but artifacts and values are predicted reliably. For our RE model, we observe that value-parameter relations are more reliably predicted than parameter-artifact relations.
@@ 5.2 LLMs
The models are run as follows. GPT-3.5 is accessed through its official API. All open models are run on a high performance compute cluster. Vicuna and WizardLM are run on nodes with NVIDIA Tesla V100 (32 GB). GALACTICA, Falcon, and Vicuna are run on nodes with NVIDIA A100 (80 GB).
@@ 5.2.2 Analysis
In Figure 6 we show an analysis of the steps leading up to model prediction. Focussing first on the zero-shot models (upper five) we observe the following across the four plots from left to right. (a) For three of five models, prompting for YAML leads to fewer parsing errors. (b) Unwanted text around the extracted data is generated more/less by two models each. (c) Hallucinated entities and (d) out of scope entities appear overall slightly more often for in YAML compared to JSON. For our few-shot approach (bottom model), we see that the use of a grammar (a, b) prevents all output format issues. Furthermore (c) hallucinated entities are reduced. (d) Out of scope entities can not be evaluated, because our in-context examples lead to frequent omission of type information in the output.
@@ 6 Discussion
Our overall results, with a top performance of 79% for entity recognition and 39% for relation extraction, show that extraction of hyperparameter information from scientific text can be accomplished to a degree that yields sound results. There are, however, challenges that remain, such as more reliable entity recognition of parameters and contexts, as well as more reliable relation extraction in general. Our novel data set enables further development of approaches from hereon. Our IE results on large-scale unannotated data give an indication of possible downstream analyses and applications. Here we see large potential for reproducibility research, faceted search, and recommendation.
Our LLM evaluation shows that for IE tasks dealing with complex information, the choice of text based data serialization format can have a considerable impact on performance, even when using grammar based generation constrains. Additionally, we can see that in-context learning enabled by larger context sizes, as well as grammars, are an effective method to improve IE performance.
6.0.1 Limitations
(1) Our work considers HyperPIE from text. This is sensible for a focussed approach, but downstream applications could furthermore benefit from composite pipelines also targeting extraction from tables, source code, etc. (2) We do not test transferability of methods to domains outside of ML related fields. It would require domain expertise to find useful definitions for hyperparameters in each respective domain. (3) Our LLM evaluation does not cover fine-tuning. Presupposing the existence of a large enough training data set, this would be a valuable addition the overall investigation. (4) Defining our YAML/JSON output format hierarchically means that only values associated with parameters and parameters associated with artifacts can be extracted. (5) Lastly, our data and experiments unfortunately are limited to English text only and do not cover other languages.